# Stochastic Calculus for Pathwise Observables of Markov-Jump Processes: Unification of Diffusion and Jump Dynamics

### v1

@@ V.1 Derivation of Response Formula for MJP
$\langle O(t)\rangle_{(\delta)}=\sum_{(x_{\tau})_{0\leq\tau\leq t}}O[(x_{\tau})_{0\leq\tau\leq t}]\mathds{P}^{(\delta)}[(x_{\tau})_{0\leq\tau\leq t}].$ (134)
@@
$\mathds{P}^{\delta}[(x_{\tau})_{0\leq\tau\leq t}]=p_{x_{0}}(0)\exp\left[\int_{s=0}^{s=t}\sum_{x}r^{\delta}_{xx}\mathrm{d}\tau_{x}(s)+\int_{s=0}^{s=t}\sum_{x,y\neq x}\log r_{xy}^{\delta}\mathrm{d}n_{xy}(s)\right],$ (135)
where $p_{x_{0}}(0)$ is the distribution of the initial state $x_{\tau=0}=x_{0}$. Note that a given path realization defines a unique sequence of transitions and hence $\mathrm{d}\tau_{x}(s)$ and $\mathrm{d}n_{xy}(s)$, which in turn enter equally $\mathds{P}$ in Eq. (33) and $\mathds{P}^{\delta}$ in Eq. (135). However, because of the different rates ($r^{\delta}_{xy}$ versus $r_{xy}$), the measure of the same path is generally different under $\mathds{P}$ and $\mathds{P}^{\delta}$. Moreover, writing the path measure in terms of stochastic differentials as in Eq. (135) further has the advantage of allowing for time-dependent rates.
@@
Note that we assume that the rates $r_{xy}=r_{xy}(\chi)$ and $r_{xy}^{\delta}$ are functions of $\chi$ and $\chi+\delta$, respectively, ($\chi$ is a placeholder for the perturbed quantity, e.g., temperature $T$) and that the function is analytic at $\delta=0$, i.e., that it has a well-defined Taylor expansion at that point. Therefore, $\mathds{P}^{\delta}$ is absolutely continuous w.r.t. $\mathds{P}$ for sufficiently small $\delta$  [^3] For MJP, the perturbed measure $\mathds{P}^{\delta}\ll\mathds{P}$ is absolutely continuous w.r.t. the unperturbed measure $\mathds{P}$ as long as for all $x\neq y$ $r_{xy}>0\implies r_{xy}^{\delta}>0$, i.e., as long as the underlying topology of the MJP remains unchanged. and the perturbed average $\langle O(t)\rangle_{\delta}$ can be written in terms of the unperturbed process in Eq. (39) using the Cameron-Martin-Girsanov theorem with the Radon-Nikodym derivative [163]
$\langle O(t)\rangle_{\delta}=\left\langle O(t)\frac{\mathrm{d}\mathds{P}^{\delta}}{\mathrm{d}\mathds{P}}\right\rangle$ (137)
@@
$\frac{\mathrm{d}\mathds{P}^{\delta}}{\mathrm{d}\mathds{P}}\left[(x_{\tau})_{0\leq\tau\leq t}\right]$ $=\exp\left[\int_{s=0}^{s=t}\sum_{x}(r_{xx}^{\delta}-r_{xx})\mathrm{d}\tau_{x}(s)+\int_{s=0}^{s=t}\sum_{x,y\neq x}\log\frac{r_{xy}^{\delta}}{r_{xy}}\mathrm{d}n_{xy}(s)\right]$
@@ V.2 Equilibrium Response to Temperature Perturbations
The aim here is to relate the response in Eq. (139) of a single-time observable $A(t)$ to the equilibrium correlation function,
@@
V.3 Examples
@@
For simplicity we set $D=T$ in Eq. (140). In Fig. 11a and Fig. 11c we use Eq. (63) with the free energies listed in Tab. 4. In Fig. 11a, we compare $\lambda=0.3$ and $\lambda=0.9$, while in Fig. 11c we use $\lambda=0.5$. The rates used in Fig. 11b and Fig. 11d are $r_{12}=r_{23}=r_{34}=r_{41}=T\mathrm{e}^{-\log\epsilon/T}$ in one direction and $r_{21}=r_{32}=r_{43}=r_{14}=T\mathrm{e}^{-\log\alpha/T}$ in the opposite direction with $\epsilon=3$ and $\alpha=2$, respectively. The temperature is set to $T=1$ and equilibrium initial conditions $p^{\mathrm{init}}_{i}=p^{\mathrm{eq}}_{i}\propto e^{-E_{i}/T}$ are assumed in Fig. 11a, while $T=2$ and initial condition $p^{\mathrm{init}}=(0,0,1,0)$ is used in Fig. 11b-Fig. 11d. The diamonds on the x-axis denote the largest relaxation time of the respective generators.
@@ VII Time-Dependent Driving
$\mathrm{cov}({J}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[\mathds{1}_{\tau>\tau^{\prime}}r_{xy}P(x,\tau|i,\tau^{\prime};v)p_{i}(\tau^{\prime};v)+\mathds{1}_{\tau<\tau^{\prime}}r_{yx}(v\tau)P(i,\tau^{\prime}|y,\tau;v)p_{y}(\tau;v)-r_{xy}(v\tau)p_{x}(\tau;v)p_{i}(\tau^{\prime};v)\right],$
@@
$+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[\mathds{1}_{\tau>\tau^{\prime}}r_{xy}(v\tau)r_{ji}(v\tau^{\prime})P(x,\tau|j,\tau^{\prime};v)p_{i}(\tau^{\prime};v)+\mathds{1}_{\tau<\tau^{\prime}}r_{ij}(v\tau^{\prime})r_{yx}(v\tau)P(i,\tau^{\prime}|y,\tau;v)p_{y}(\tau;v)\right$.
@@
$\langle A_{t}\rho_{t}\rangle=$ $\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i}V_{i}(v\tau^{\prime})\mathds{1}_{\tau<\tau^{\prime}}\sum_{x,y}\left[P(i,\tau^{\prime}|y,\tau;v)-P(i,\tau^{\prime}|x,\tau;v)\right]p_{x}(\tau;v)r_{xy}(v\tau)Z_{xy}(v\tau).$ (159)
@@
$\langle$ $A_{t}\rho_{t}\rangle$ (160)
$=$ $\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{\tau^{\prime}}\mathrm{d}\tau\sum_{i,x}V_{i}(v\tau^{\prime})P(i,\tau^{\prime}|x,\tau;v)\partial_{\tau}p_{x}(\tau;v)$
$=$ $-\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{\tau^{\prime}}\mathrm{d}\tau\sum_{i,x}V_{i}(v\tau^{\prime})\left[\partial_{\tau}P(i,\tau^{\prime}|x,\tau;v)\right]p_{x}(\tau;v),$
@@
VIII Outlook

### v2

@@ V.1 Derivation of Response Formula for MJP
$\langle O(t)\rangle_{(\delta)}=\sum_{(x_{\tau})_{0\leq\tau\leq t}}O[(x_{\tau})_{0\leq\tau\leq t}]{\color[rgb]{0,0,0}\mathbb{P}}^{(\delta)}[(x_{\tau})_{0\leq\tau\leq t}].$ (142)
@@
${\color[rgb]{0,0,0}\mathbb{P}}^{\delta}[(x_{\tau})_{0\leq\tau\leq t}]=p_{x_{0}}(0)\exp\left[\int_{s=0}^{s=t}\sum_{x}r^{\delta}_{xx}\mathrm{d}\tau_{x}(s)+\int_{s=0}^{s=t}\sum_{x,y\neq x}\log r_{xy}^{\delta}\mathrm{d}n_{xy}(s)\right],$ (143)
[…]
@@
[…]
$\langle O(t)\rangle_{\delta}=\left\langle O(t)\frac{\mathrm{d}{\color[rgb]{0,0,0}\mathbb{P}}^{\delta}}{\mathrm{d}{\color[rgb]{0,0,0}\mathbb{P}}}\right\rangle$ (145)
@@
$\frac{\mathrm{d}{\color[rgb]{0,0,0}\mathbb{P}}^{\delta}}{\mathrm{d}{\color[rgb]{0,0,0}\mathbb{P}}}\left[(x_{\tau})_{0\leq\tau\leq t}\right]$ $=\exp\left[\int_{s=0}^{s=t}\sum_{x}(r_{xx}^{\delta}-r_{xx})\mathrm{d}\tau_{x}(s)+\int_{s=0}^{s=t}\sum_{x,y\neq x}\log\frac{r_{xy}^{\delta}}{r_{xy}}\mathrm{d}n_{xy}(s)\right]$
@@ V.2 Equilibrium Response to Temperature Perturbations
The aim here is to relate the response in Eq. (147) of a single-time observable $O(t)$ to the equilibrium correlation function,
@@
V.3 Connection to Response Function Formalism
In this section, we connect the results derived in the previous sections to the well established response function formalism, which is often used to describe the fluctuation dissipation theorem. While the results of response functions for MJP has been previously derived in, e.g., [64, 110, 153, 173, 142] , we show how these may be extracted from our results to bridge the “gap” to the literature.
V.3.1 Equilibrium Response
For a first order perturbation theory of a single-time observable $B(t)$ in an equilibrium system with perturbation $h(s)$, $0\leq s\leq t$, conjugate to some observable $R(s)$, the response reads [64, 110]
$\Delta B(t)\equiv\langle B(t)\rangle_{h}-\langle B\rangle_{\mathrm{eq}}=\int_{0}^{t}\mathrm{d}sh(s)\chi(t-s)\,.$ (155)
Here, $\chi(s)$ is the response function (recall that $k_{\mathrm{B}}=1$ throughout the manuscript) [64]
$\chi(s)=-\frac{1}{T}\frac{\mathrm{d}}{\mathrm{d}s}\langle R(0)B(s)\rangle_{\mathrm{eq}}\,.$ (156)
To connect this to the result we derived in the previous section, we first consider the case a constant perturbation to an equilibrium system, i.e., $h(s)=h$. Comparing Eq. (154) to Eq. (155) (with $\delta\to h$), we see that to first order in $h$
$\Delta O(t)$ $=\frac{h}{T^{2}}\left(\langle OE\rangle_{\mathrm{eq}}-\langle O(t)E(0)\rangle_{\mathrm{eq}}\right)$
$=\frac{1}{T^{2}}\int_{0}^{t}\mathrm{d}s\frac{\mathrm{d}}{\mathrm{d}s}\langle O(t)E(s)\rangle_{\mathrm{eq}}h\,.$ (157)
Since the equilibrium correlation is time-translation invariant, we can rewrite the above response Eq. (157) as
$\Delta O(t)=\int_{0}^{t}\mathrm{d}s\chi(t-s)h+\mathcal{O}(h^{2})\,$, (158)
with $\chi(t-s)\equiv\frac{1}{T^{2}}\frac{\mathrm{d}}{\mathrm{d}s}\langle{O}(t)E(s)\rangle_{\mathrm{eq}}=-\frac{1}{T^{2}}\frac{\mathrm{d}}{\mathrm{d}t}\langle O(t-s)E(0)\rangle_{\mathrm{eq}}$ consistent with Eq. (156) with $R=T^{-1}E$.
For time-dependent $h(s)$ we get to first order
$\Delta O(t)=\left\langle O(t)\int_{s=0}^{s=t}h(s)\sum_{x\neq y}\frac{\tilde{r}_{xy}}{r_{xy}}\mathrm{d}\varepsilon_{xy}(s)\right\rangle\,.$ (159)
The simplifications performed in Eqs. (151), (152), and (153) can still be made, as these are independent of the perturbation $h(s)$
$\Delta O(t)$ $=\frac{1}{T^{2}}\int_{0}^{t}h(s)\sum_{i}\sum_{x}O_{i}E_{x}\partial_{s}P(i,t|x,s)p_{x}^{\mathrm{eq}}\mathrm{d}s$
$=\frac{1}{T^{2}}\int_{0}^{t}h(s)\partial_{s}\langle O(t)E(s)\rangle_{\mathrm{eq}}\mathrm{d}s$
$=\int_{0}^{t}h(s)\chi(t-s)\,$, (160)
again with
$\chi(t-s)$ $=\frac{1}{T^{2}}\frac{\mathrm{d}}{\mathrm{d}s}\langle O(t)E(s)\rangle_{\mathrm{eq}}$
$=-\frac{1}{T^{2}}\frac{\mathrm{d}}{\mathrm{d}t}\langle O(t-s)E(0)\rangle_{\mathrm{eq}}\,.$ (161)
The extension of these results, e.g., in case of non-stationary systems, has already been discussed in Ref. [64].
V.3.2 Response of a Stationary System
To further connect our result to responses of (non-equilibrium) steady states, we realize that the perturbation rates $\tilde{r}_{xy}$ are the elements of the perturbation generator $\bm{L}^{1}$, i.e., $\bm{L}=\bm{L}_{0}+h\bm{L}^{1}$ with $L_{xy}=r_{yx}$ and $L^{1}_{xy}=\tilde{r}_{yx}$. Using the correlation lemma Eq. (50), the response of a single-time observable in steady-state reads to first order in $h$
$\Delta_{\mathrm{s}}O(t)\equiv$ $\langle O(t)\rangle_{h}-\langle O(t)\rangle_{\mathrm{s}}$
$=$ $\int_{0}^{t}\mathrm{d}sh(s)\sum_{i}O_{i}\sum_{x,y\neq x}\frac{\tilde{r}_{xy}}{r_{xy}}$
$\times\left[P(i,t|y,s)-P(i,t|x,s)\right]r_{xy}p_{x}^{\mathrm{s}}\,.$ (162)
The last double-sum simplifies as follows
$\sum_{y,x\neq y}\tilde{r}_{xy}[P(i,t|y,s)-P(i,t|x,s)]p_{x}^{\mathrm{s}}$
$=$ $\sum_{y,x\neq y}\tilde{r}_{yx}P(i,t|x,s)p_{y}^{\mathrm{s}}-\underbrace{\sum_{y,x\neq y}\tilde{r}_{xy}}_{\tilde{r}_{xx}}P(i,t|x,s)]p_{x}^{\mathrm{s}}$
$=$ $\sum_{y,x}\tilde{r}_{yx}P(i,t|x,s)p_{y}^{\mathrm{s}}\,$, (163)
where we split the sum and relabel $x\leftrightarrow y$ in the first line, and then use the definition of the diagonal elements to simplify the expression. As a consequence, we can now write the conjugate observable $R_{x}^{\mathrm{s}}\equiv\sum_{y}L_{xy}^{1}p_{y}^{\mathrm{s}}/p_{x}^{\mathrm{s}}$, and identify that inserting Eq. (163) into Eq. (162) yields to first order
$\Delta_{\mathrm{s}}O(t)=$ $\int_{0}^{t}\mathrm{d}sh(s)\sum_{i}O_{i}\sum_{y,x}\underbrace{\tilde{r}_{yx}}_{L_{xy}^{1}}P(i,t|x,s)p_{y}^{\mathrm{s}}$
$=$ $\int_{0}^{t}\mathrm{d}sh(s)\sum_{i}O_{i}\sum_{x}P(i,t|x,s)\underbrace{\sum_{y}L_{xy}^{1}\frac{p_{y}^{\mathrm{s}}}{p_{x}^{\mathrm{s}}}}_{R^{\mathrm{s}}_{x}}p_{x}^{\mathrm{s}}$
$=$ $\int_{0}^{t}\mathrm{d}sh(s)\sum_{i,x}O_{i}P(i,t|x,s)R^{\mathrm{s}}_{x}p_{x}^{\mathrm{s}}\,.$ (164)
Here, we can identify the stationary response function
$\chi_{\mathrm{s}}(t-s)\equiv{\color[rgb]{0,0,0}\mathbb{1}}_{s<t}\langle O(t)R^{\mathrm{s}}(s)\rangle_{\mathrm{s}}={\color[rgb]{0,0,0}\mathbb{1}}_{s<t}\langle O(t-s)R^{\mathrm{s}}(0)\rangle_{\mathrm{s}}\,$, (165)
so that Eq. (164) can be written as
$\Delta_{\mathrm{s}}O(t)=\int_{0}^{t}\mathrm{d}sh(s)\chi_{\mathrm{s}}(t-s)+\mathcal{O}(h^{2})\,.$ (166)
This is consistent with the results from Ref. [142].
V.3.3 Response of Path-Wise Observables
Lastly, we consider the response of, possibly transient, path-wise observables. Specifically, consider the observables
$O_{1}(t)$ $=\int_{0}^{t}\sum_{x,y\neq x}b_{xy}(s)\mathrm{d}\varepsilon_{xy}(s)\,$, (167)
$O_{2}(t)$ $=\int_{0}^{t}g_{s}\mathrm{d}s\,.$ (168)
Using Eq. (48) the response of Eq. (167) can be stated directly to first order
$\Delta O_{1}(t)\equiv$ $\langle O_{1}(t)\rangle_{h}-\langle O_{1}(t)\rangle$
$=$ $\int_{0}^{t}\mathrm{d}z\sum_{x,y\neq x}b_{xy}(z)\int_{0}^{t}\mathrm{d}sh(s)\sum_{i,j\neq i}\frac{\tilde{r}_{ij}}{r_{ij}}$
$\times\delta_{ix}\delta_{jy}\delta(s-z)r_{ij}p_{i}(s)$
$=$ $\int_{0}^{t}\mathrm{d}z\sum_{x,y\neq x}b_{xy}(z)h(z)\tilde{r}_{xy}p_{x}(z)\,.$ (169)
Similarly, using Eq. (50), the response of $O_{2}(t)$ becomes to first order (note that this result is compared to the approach used in Ref. [64] in App. A.8)
$\Delta O_{2}(t)\equiv$ $\langle O_{2}(t)\rangle_{h}-\langle O_{2}(t)\rangle$
$=$ $\int_{0}^{t}\mathrm{d}z\int_{0}^{t}\mathrm{d}s{\color[rgb]{0,0,0}\mathbb{1}}_{s<z}h(s)\sum_{i,x,y\neq x}$ (170)
$\times g_{i}(z)\tilde{r}_{xy}p_{x}(s)\left[P(i,z|y,s)-P(i,z|x,s)\right].$
The sum over $x,y$ can again be simplified using Eq. (163), so that
$\Delta O_{2}(t)=$ $\int_{0}^{t}\mathrm{d}z\int_{0}^{t}\mathrm{d}s{\color[rgb]{0,0,0}\mathbb{1}}_{s<z}h(s)\sum_{i,x,y}g_{i}(z)\tilde{r}_{yx}p_{y}(s)P(i,z|x,s)$
$+\mathcal{O}(h^{2})\,.$ (171)
Defining $R_{x}(s)=\sum_{y}\tilde{r}_{yx}p_{y}(s)/p_{x}(s)=\sum_{y}L_{xy}^{1}p_{y}(s)/p_{x}(s)$}, the response functions for $O_{1}(t)$ and $O_{2}(t)$ become
$\chi_{O_{1}}(t,s)$ $={\color[rgb]{0,0,0}\mathbb{1}}_{t>s}\sum_{x,y\neq x}b_{xy}(s)\tilde{r}_{xy}p_{x}(s),$ (172)
$\chi_{O_{2}}(t,s)$ $=\int_{0}^{t}\mathrm{d}z{\color[rgb]{0,0,0}\mathbb{1}}_{z>s}\langle g_{z}R(s)\rangle,$ (173)
so that the response to any additive functional $O_{\mathrm{p}}(t)=O_{1}(t)+O_{2}(t)$ becomes
$\Delta O_{\mathrm{p}}(t)\equiv$ $\langle O_{\mathrm{p}}\rangle_{h}-\langle O_{\mathrm{p}}\rangle$
$=$ $\int_{0}^{t}\mathrm{d}sh(s)\chi_{O_{\mathrm{p}}}(t,s)+\mathcal{O}(h^{2})\,$,
$\chi_{O_{\mathrm{p}}}(t,s)=$ $\chi_{O_{1}}(t,s)+\chi_{O_{2}}(t,s)\,.$ (174)
Note that observables of integrals over $\mathrm{d}\bm{n}(s)$ is a special case where $g_{x}(s)$ and $b_{xy}(s)$ have been chosen appropriately to satisfy Eq. (46).
V.4 Examples
@@
For simplicity we set $D=T$ in Eq. (148). In Fig. 11a and Fig. 11c we use Eq. (71) with the free energies listed in Tab. 5. In Fig. 11a, we compare $\lambda=0.3$ and $\lambda=0.9$, while in Fig. 11c we use $\lambda=0.5$. The rates used in Fig. 11b and Fig. 11d are $r_{12}=r_{23}=r_{34}=r_{41}=T\mathrm{e}^{-\log\epsilon/T}$ in one direction and $r_{21}=r_{32}=r_{43}=r_{14}=T\mathrm{e}^{-\log\alpha/T}$ in the opposite direction with $\epsilon=3$ and $\alpha=2$, respectively. The temperature is set to $T=1$ and equilibrium initial conditions $p^{\mathrm{init}}_{i}=p^{\mathrm{eq}}_{i}\propto e^{-E_{i}/T}$ are assumed in Fig. 11a, while $T=2$ and initial condition $p^{\mathrm{init}}=(0,0,1,0)$ is used in Fig. 11b-Fig. 11d. The diamonds on the horizontal axis denote the largest relaxation time of the respective generators.
@@ VII Time-Dependent Driving
$\mathrm{cov}({J}_{t}^{k},{\rho}_{t}^{l})=$ $\hat{{I}}^{t}_{\kappa^{k},V^{l}\delta}\left[{\color[rgb]{0,0,0}\mathbb{1}}_{\tau>\tau^{\prime}}r_{xy}P(x,\tau|i,\tau^{\prime};v)p_{i}(\tau^{\prime};v)+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}r_{yx}(v\tau)P(i,\tau^{\prime}|y,\tau;v)p_{y}(\tau;v)-r_{xy}(v\tau)p_{x}(\tau;v)p_{i}(\tau^{\prime};v)\right],$
@@
$+\hat{{I}}^{t}_{\kappa^{k},\kappa^{l}}\left[{\color[rgb]{0,0,0}\mathbb{1}}_{\tau>\tau^{\prime}}r_{xy}(v\tau)r_{ji}(v\tau^{\prime})P(x,\tau|j,\tau^{\prime};v)p_{i}(\tau^{\prime};v)+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}r_{ij}(v\tau^{\prime})r_{yx}(v\tau)P(i,\tau^{\prime}|y,\tau;v)p_{y}(\tau;v)\right$.
@@
$\langle A_{t}\rho_{t}\rangle=$ $\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i}V_{i}(v\tau^{\prime}){\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\sum_{x,y}\left[P(i,\tau^{\prime}|y,\tau;v)-P(i,\tau^{\prime}|x,\tau;v)\right]p_{x}(\tau;v)r_{xy}(v\tau)Z_{xy}(v\tau).$ (187)
@@
$\langle$ $A_{t}\rho_{t}\rangle\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{\tau^{\prime}}\mathrm{d}\tau\sum_{i,x}V_{i}(v\tau^{\prime})P(i,\tau^{\prime}|x,\tau;v)\partial_{\tau}p_{x}(\tau;v)=-\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{\tau^{\prime}}\mathrm{d}\tau\sum_{i,x}V_{i}(v\tau^{\prime})\left[\partial_{\tau}P(i,\tau^{\prime}|x,\tau;v)\right]p_{x}(\tau;v),$
@@
VIII Connection to Quantum Unraveling
Stochastic calculus has also found traction in the theory of open quantum systems [13, 50, 25, 27, 9], i.e., in the context of quantum unraveling. In the following, we provide some insight into how the classical stochastic-calculus approach presented here is connected to quantum unraveling. Before doing so, we introduce some basic notation and theory. Subsequently, we highlight the relation to our work.
Let $\rho_{s}^{\mathrm{u}}$ be a rank-1 matrix denoting a pure state in a discrete-state open quantum system at time $s$. It evolves according to the Belavkin equation [27]
$\mathrm{d}\rho_{s}^{\mathrm{u}}=\mathcal{B}(\rho_{s}^{\mathrm{u}})\mathrm{d}s+\sum_{i}\left(\frac{\mathcal{J}_{i}(\rho_{s}^{\mathrm{u}})}{\mathrm{Tr}[\mathcal{J}_{i}(\rho_{s}^{\mathrm{u}})]}-\rho_{s}^{\mathrm{u}}\right)\mathrm{d}\tilde{n}_{is},$ (191)
where $\mathcal{J}_{i}(\rho_{s}^{\mathrm{u}})=\hat{J}_{i}\rho_{s}^{\mathrm{u}}\hat{J}_{i}^{\dagger}$ with jump operator $\hat{J}_{i}$ for the $i$th quantum jump. The first term in Eq. (191) has the form
$\mathcal{B}(\rho_{s}^{\mathrm{u}})=-i\hat{H}_{\mathrm{eff}}\rho_{s}^{\mathrm{u}}+\rho_{s}^{\mathrm{u}}\hat{H}_{\mathrm{eff}}^{\dagger}-\rho_{s}^{\mathrm{u}}\mathrm{Tr}(-i\hat{H}_{\mathrm{eff}}\rho_{s}^{\mathrm{u}}+\rho_{s}^{\mathrm{u}}\hat{H}_{\mathrm{eff}}^{\dagger}),$ (192)
where the Hamiltonian $\hat{H}$ of the system enters through the effective Hamiltonian $\hat{H}_{\mathrm{eff}}=\hat{H}-\frac{i}{2}\sum_{j}\hat{J}_{j}^{\dagger}\hat{J}_{j}$. Lastly, $\mathrm{d}\tilde{n}_{is}$ is a stochastic differential which takes values 1 if the $i$th transition occurs in $s,s+\mathrm{d}s]$ and 0 otherwise. It has the properties $\langle\mathrm{d}\tilde{n}_{is}\rangle=\mathrm{Tr}(\mathcal{J}_{i}(\rho_{s}))$ and $\mathrm{d}\tilde{n}_{is}\mathrm{d}\tilde{n}_{js}=\delta_{ij}\mathrm{d}\tilde{n}_{is}$. Taking the average over noise histories in Eq. (191), the resulting density matrix $\rho_{s}=\langle\rho_{s}^{\mathrm{u}}\rangle$ evolves according to the Lindblad equation, often referred to as “quantum master equation”.
With this short insight into quantum SDE’s, we now turn to the question of how these concepts are related to the work we present in this article. Since classical stochastic dynamics should correspond to a special case of quantum open-system evolutions [22], one may at first think that Eq. (46) describes the classical counterpart to Eq. (191); they both aim to describe a trajectory (the former a classical sequence of discrete states and the latter the evolution of pure states in a Hilbert space) and both include Poissonian noise $\mathrm{d}\bm{n}$ and $\mathrm{d}\tilde{n}$, respectively.
However, Eq. (46) is not the classical counterpart to Eq. (191), which can be seen from two observations: (i) Eq. (46) is an SDE describing the evolution of transitions and (ii) Eq. (191) is a functional of the transitions describing the evolution of the states. Moreover, Eq. (46) allows us to consider the (stochastic) evolution of transitions occuring in a discrete systems and further enables us to form and study additive functionals of the dynamics. Hence, using (i) enables us to study functionals such as (ii). To be precise on this last point, we can construct a classical functional of the transitions which is the classical counterpart to Eq. (191). It turns out that this functional simply is $\mathrm{d}{\color[rgb]{0,0,0}\delta_{x_{s}x}}\equiv\frac{\mathrm{d}}{\mathrm{d}\tau}{\color[rgb]{0,0,0}\delta_{x_{s}x}}\mathrm{d}\tau$, see Eq. (212). Explicitly, using Eq. (46) in Eq. (212) yields
$\mathrm{d}{\color[rgb]{0,0,0}\delta_{x_{s}x}}$ $=\sum_{y\neq x}(\mathrm{d}n_{yx}-\mathrm{d}n_{xy})$ (193)
$=\sum_{y\neq x}(r_{yx}\delta_{x_{s}y}\mathrm{d}s-r_{xy}\delta_{x_{s}x}\mathrm{d}s)+\sum_{y\neq x}(\mathrm{d}\varepsilon_{yx}-\mathrm{d}\varepsilon_{xy}).$
The first sum is the deterministic evolution of the state, i.e., the master operator acting on the “population vector” $\delta_{x_{s}x}$ and the second sum is the zero-mean stochastic noise in the system. Shifting the mean of the noise increments by adding and subtracting $\sum_{y\neq x}(\langle\mathrm{d}n_{yx}\rangle-\langle\mathrm{d}n_{xy}\rangle)$ [this corresponds to the trace terms in Eq. (191)], we recover the classical analogue of Eq. (191)
$\mathrm{d}\delta_{x_{s}x}=$ $\sum_{y\neq x}\left(r_{yx}\delta_{x_{s}y}\mathrm{d}s-\langle\mathrm{d}n_{yx}\rangle-r_{xy}\delta_{x_{s}x}\mathrm{d}s+\langle\mathrm{d}n_{xy}\rangle\right)$
$+\sum_{y\neq x}(\mathrm{d}\varepsilon_{yx}+\langle\mathrm{d}n_{yx}\rangle-\mathrm{d}\varepsilon_{xy}-\langle\mathrm{d}n_{xy}\rangle).$ (194)
Indeed, taking the average of Eq. (193) w.r.t. the noise history recovers the master equation, i.e., the classical analogue of the Lindblad equation, further showing the correspondence between Eqs. (191) and (193).
For completeness it should be mentioned that the quantum analogue of Eq. (46) has already been identified in Ref. [93] as
$\mathrm{d}\tilde{n}_{it}=\mathrm{Tr}(\mathcal{J}_{i}(\rho_{t}))\mathrm{d}t+\mathrm{d}\tilde{\varepsilon}_{it},$ (195)
where the noise $\mathrm{d}\tilde{\varepsilon}_{it}$ corresponds to Eq. (45) and the statistical properties can be evaluated in a similar fashion; $\langle\mathrm{d}\tilde{\varepsilon}_{it}\rangle=0$ and $\langle\mathrm{d}\tilde{\varepsilon}_{it}\mathrm{d}\tilde{\varepsilon}_{jt^{\prime}}\rangle=\delta_{ij}\delta(t-t^{\prime})\mathrm{Tr}(\mathcal{J}_{i}(\rho_{t}))\mathrm{d}t\mathrm{d}t^{\prime}$. While the connection to the unraveling of the Lindblad equation and, thus, the Belavkin equation is therefore established, the question of how this in turn relates to the classical counterpart remains elusive, further motivating our discussion here.
IX Outlook